OpenStudy (anonymous):

what is the inverse of f(x)=(x+3)/(x-2)

OpenStudy (anonymous):

Switch Xs and Ys

OpenStudy (anonymous):

Then solve for y

OpenStudy (anonymous):

i just want to check because my math books syas the anser is (2x+3)/(x-1)

OpenStudy (anonymous):

but, on the last problem, my book had a mistake so now i dont trust it

myininaya (myininaya):

\[y=\frac{x+3}{x-2}\] \[y(x-2)=x+3\] \[yx-2y=x+3\] \[yx-x=3+2y\] \[x(y-1)=3+2y\] \[\frac{x(y-1)}{y-1}=\frac{3+2y}{y-1}\] \[x=\frac{3+2y}{y-1}\] now just interchange y with x and x with f^(-1)(x)

OpenStudy (anonymous):

could you also check what the inverse for this one is f(x)=x^3+1. My books says the anser is \[5\sqrt{x-1}\]

myininaya (myininaya):

just solve for x and replace x with f^(-1)(x) and replace y with x let me see your attempt

OpenStudy (anonymous):

well i got and everyone else is telling me that the answer is \[\sqrt[3]{x-1}\]

myininaya (myininaya):

you have \[y=x^3+1\] how are we gonna solve this for x?

OpenStudy (anonymous):

we first subtract 1 from both sides, then cub root both sides

myininaya (myininaya):

right so after subtracting one on both sides we get: \[y-1=x^3\] then take cube root of both sides \[\sqrt[3]{y-1}=x\] now replace y with x and x with f^(-1)(x)

OpenStudy (anonymous):

alright, guess my book has a mistake

myininaya (myininaya):

yes